Logiweb bibliography of check

Up Help

[0] check

[1] base

The pyk compiler, version 0.grue.20050502+ by Klaus Grue GRD-2005-05-26.UTC:06:49:32.964854 = MJD-53516.TAI:06:50:04.964854 = LGT-4623807004964854e-6